# Spatial Context-Driven Positive Pair Sampling for Enhanced Histopathology Image Classification

### v1

@@ Abstract
Deep learning has demonstrated great promise in cancer classification from whole-slide images (WSIs) but remains constrained by the need for extensive annotations. Annotation-free methods, such as multiple instance learning (MIL) and self-supervised learning (SSL), have emerged to address this challenge; however, current SSL techniques often depend on synthetic augmentations or temporal context, which may not adequately capture the intricate spatial relationships inherent to histopathology. In this work, we introduce a novel spatial context-driven positive pair sampling strategy for SSL that leverages the natural coherence of adjacent patches in WSIs. By constructing biologically relevant positive pairs from spatially proximate patches, our approach harnesses inherent spatial coherence to enhance patch-level representations, ultimately boosting slide-level classification performance. Experiments on multiple datasets reveal that our strategy improves classification accuracy by 5% to 10% over the standard method, paving the way for more clinically relevant AI models in cancer diagnosis. The code is available at https://anonymous.4open.science/r/contextual-pairs-E72F/.
Keywords:
Cancer Diagnosis Digital Pathology Whole-Slide Images Self-Supervised Learning Positive Pair Sampling.
@@ 1 Introduction
Deep learning (DL) has transformed cancer classification with whole-slide images (WSIs), achieving pathologist-level performance [15, 19, 20]. However, training these models requires large labeled datasets, which are scarce and labor-intensive to annotate, as they depend on expert pathologists [17, 25].
To address the scarcity of labeled data, annotation-free approaches have been proposed [14, 16, 23, 26]. One prominent method is Multiple Instance Learning (MIL), which treats each WSI as a "bag" of image patches with a single slide-level label [7]. Another promising strategy is self-supervised learning (SSL), which enables models to learn from vast amounts of unlabeled data. SSL has proven particularly valuable in histopathology, where obtaining expert-annotated datasets is costly and time-consuming. Notably, SSL pre-training on pathology data consistently outperforms ImageNet pre-training, leading to better representations for downstream pathology tasks [15, 19, 20].
While SSL enables feature learning without labeled data [15, 18], adapting it to pathology images presents unique challenges. Unlike natural images, pathology images lack a canonical orientation and exhibit low color variation [15]. Standard SSL techniques, such as random flips, stain augmentations, and multi-view training, attempt to address these issues [8, 24, 6, 15]. However, these synthetic transformations may not fully capture the natural variations found in real-world tissue samples.
Recent work in natural image domains suggests that leveraging spatial or temporal relationships within unlabeled data can generate more meaningful positive pairs for SSL [21, 13]. For example, Pantazis et al. [21] utilized spatial and temporal proximity to create informative training pairs in wildlife camera images, while Tile2Vec [13] employed geographic proximity with a triplet loss to differentiate between similar and dissimilar regions in satellite images. Inspired by these findings, we propose a novel SSL approach that incorporates spatial proximity for more biologically relevant positive pair sampling in WSI analysis.
@@
Figure 1: In the standard approach (blue box), a patch from the WSI is transformed to create a standard positive pair for self-supervised learning. Our method (green box) additionally leverages spatial proximity to sample a near-neighbor patch, forming a contextual positive pair. By combining both pairs, we exploit transformation invariance and contextual similarity, resulting in more robust representation learning.
@@
In this paper, we introduce and evaluate our contextual positive pair sampling strategy within the SSL framework (illustrated in Figure 1). Rather than aiming to surpass state-of-the-art (SOTA) performance, our goal is to present a biologically grounded method that enhances representation learning for histopathology, aligning with the growing interest in foundational models for medical imaging [15, 20, 19]. By leveraging spatial coherence between adjacent image patches, our approach constructs more meaningful positive pairs, leading to improved WSI classification performance over standard SSL sampling strategies.
@@ 2 Methodology
2.1 Self-supervised Learning
Self-supervised Learning (SSL) methods aim to train an embedding function that maps an input image $x$ to a meaningful representation, enabling downstream tasks without requiring extensive labeled data. Recent SSL methods adopt a joint embedding architecture, where two networks are optimized to produce similar embeddings for different augmented views of the same image [3, 10, 28, 1, 5].
In the context of WSI analysis, an input image $x$ is a small patch extracted from a WSI, denoted as $P=\{p_{1},...,p_{m}\}$. Given a set of image transformation $T$, transformations $t_{1},t_{2}\sim T$ are sampled to two views: $v_{1,i}=t_{1}(p_{i})$ and $v_{2,i}=t_{2}(p_{i})$. An encoder $f$ is used to produce representations $z_{1}=f(v_{1,i})$ and $z_{2}=f(v_{2,i})$. A projector $h$ is commonly used to produce projections $y_{1}=h(z_{1})$ and $y_{2}=h(z_{2})$. A loss function $L(y_{1},y_{2})$ is then minimized to enforce similarity in the embedding space. Subsequently, the learned patch-level embeddings are aggregated into a slide-level representation through a features aggregator. Finally, this slide-level representation is utilized by a dedicated slide-level model for downstream classification tasks, such as cancer diagnosis.
Although earlier SSL methods like SimCLR [6] rely on negative pairs, recent approaches opt to not use negative pairs primarily for computational efficiency and improved scalability [10, 28, 1, 5]. Moreover, in benign WSIs, all extracted patches originate from benign tissue, making it infeasible to construct true negative pairs. Consequently, our method focuses exclusively on positive pairs.
2.2 Contextual Positive Pairs Sampling
The overview of our method is illustrated in Fig 1. Our framework leverages spatial proximity to construct positive pairs that capture shared underlying characteristics. Before training the SSL model, we sample $M$ patches, termed pivot patches, from all WSIs to form the training set $U^{P}$. For each pivot patch, we identify $K$ spatially proximate patches within the same WSI, referred to as neighbor patches ($U^{N}$), based on the Chebyshev distance [4] (given that WSIs are fundamentally grid-based structures). During training, a neighbor patch is randomly selected for each pivot patch, provided its distance falls within predefined thresholds, forming a contextual pair $\{p^{i}_{j}\in U^{P},p^{i}_{k}\in U^{N}\}$. In addition, we construct standard positive pairs for each pivot patch by applying two independent transformations, $t_{1}$ and $t_{2}$, yielding $\{t_{1}(p^{i}_{j}),t_{2}(p^{i}_{j})\}$. The SSL loss is computed using two components: one based on the contextual pair $(v_{1,j}=t_{1}(p^{i}_{j}),v_{1,k}=t_{1}(p^{i}_{k}))$ and another based on the standard augmentation-based pair $(v_{1,j}=t_{1}(p^{i}_{j}),v_{2,j}=t_{2}(p^{i}_{j}))$. These loss terms are weighted by a hyperparameter $\alpha$, as formulated in Equation 1:
@@ 2.2 Contextual Positive Pairs Sampling
This formulation balances contextual similarity and transformation invariance, guiding the model to learn robust representations. Moreover, it is agnostic to the specific SSL method employed, making it adaptable to various frameworks. Moreover, while various distance metrics exist, we selected Chebyshev distance because WSIs are fundamentally grid-based structures, with patches arranged in a matrix-like format.
@@ 3.1.1 Public Datasets
3.1.2 Private Datasets
We evaluated our method in two private datasets acquired from a medical foundation: (1) 600 stomach WSIs and (2) 600 colorectal WSIs. Each WSI contains 2 to 4 slices from the same tissue sample. Among the slides, and for each dataset, 200 were classified as normal/benign slides (N), 200 were classified as tissues in dysplasia (D), and 200 were classified as cancerous/malignant (M). The Institutional Review Board (Approval #SMF-IRB-2020–007) of the organization and the Institutional Review Board (Approval #KAIST-IRB-22-334, #KH2020-116) of Korea Advanced Institute of Science and Technology (KAIST) have both granted approval for research use. The WSIs were stained with hematoxylin and eosin (H&E), digitized as MIRAX files with a sensor resolution of 200x, using a Panoramic Flash 250 III scanner, and were examined and graded by two pathologists from the same organization. All patient records were anonymized. Moreover, for each class, we randomly split the WSIs into 140 slides for training, 30 slides for validation, and 30 slides for testing. Using a sliding window approach, we extracted 256x256 pixel patches at 20x magnification with the OpenSlide Python library [9].
@@ 3.2 Implementation Details
We conducted our experiments using PyTorch [22] on a Linux workstation with a single 32 GB NVIDIA GeForce RTX 3080 Ti GPU. We employed three SSL methods—Barlow Twins [28], BYOL [10], and VICReg [1]—to train the patch-level representation encoders. ResNet18 [11] was used as the backbone architecture for the SSL methods. The optimization was performed using the LARS optimizer [27], configured with a weight decay of $1\times 10^{-6}$. Separate learning rates were set for biases and weights at 0.0048 and 0.2, respectively. A batch size of 1024 was employed, and all models were trained for 100 epochs.
To generate the augmented views, we used a diverse set of transformations inspired by [6] and [15]. These augmentations included vertical and horizontal flipping, random resized cropping, color jittering, and Gaussian blur, which helped the model become invariant to spatial and visual variations.
We aggregated the patch features using Multiple Instance Learning (MIL) to make a slide-level prediction. Specifically, we used attention-based MIL (ABMIL) [12]. During MIL training, we employed the Adam optimizer with a weight decay of $1\times 10^{-2}$, a learning rate of 0.001, and a total of 100 epochs. The performance of the models was evaluated using accuracy and the Area Under the Receiver Operating Characteristic Curve (AUROC).
@@ 3.3 Contextual Positive Pair Sampling
Figure 2: Percentage of mismatches as a function of Chebyshev distance for three histopathology datasets (Stomach, Colon, and Breast).

### v2

# Leveraging Spatial Context for Positive Pair Sampling in Histopathology Image Representation Learning

@@ Abstract
Deep learning has shown strong potential in cancer classification from whole-slide images (WSIs), but the need for extensive expert annotations often limits its success. Annotation-free approaches, such as multiple instance learning (MIL) and self-supervised learning (SSL), have emerged as promising alternatives to traditional annotation-based methods. However, conventional SSL methods typically rely on synthetic data augmentations, which may fail to capture the spatial structure critical to histopathology. In this work, we propose a spatial context-driven positive pair sampling strategy that enhances SSL by leveraging the morphological coherence of spatially adjacent patches within WSIs. Our method is modular and compatible with established joint embedding SSL frameworks, including Barlow Twins, BYOL, VICReg, and DINOv2. We evaluate its effectiveness on both slide-level classification using MIL and patch-level linear probing. Experiments across four datasets demonstrate consistent performance improvements, with accuracy gains of 5% to 10% compared to standard augmentation-based sampling. These findings highlight the value of spatial context in improving representation learning for computational pathology and provide a biologically meaningful enhancement for pretraining models in annotation-limited settings. The code is available at https://anonymous.4open.science/r/contextual-pairs-E72F/.
@@ 1 Introduction
Deep learning (DL) has revolutionized cancer diagnosis using whole-slide images (WSIs), achieving pathologist-level performance in several tasks [21, 25, 27]. However, these advances rely heavily on large, expert-annotated datasets, which is an expensive and time-consuming requirement that limits scalability across institutions and cancer types [23, 41].
To mitigate the annotation bottleneck, label-efficient approaches such as Multiple Instance Learning (MIL) [9] and self-supervised learning (SSL) [19, 22, 33, 43] have gained traction. In particular, SSL has emerged as a powerful tool for learning visual representations directly from unlabeled WSI data. SSL pretraining on histopathology data has consistently outperformed ImageNet pretraining, yielding improved features for downstream classification and segmentation tasks [21, 25, 27].
Despite its promise, applying SSL to pathology introduces unique challenges. WSIs differ from natural images in that they lack consistent orientation, exhibit low color variability, and contain complex spatial tissue structures [21]. Common SSL strategies (e.g., synthetic transformations like flips or stain augmentations [10, 35, 8, 21]) may not fully capture the biological variability inherent in tissue morphology.
Recent advances in natural image domains highlight the benefit of using contextual signals, spatial or temporal, to define semantically meaningful positive pairs [30, 18]. For instance, Tile2Vec [18] uses geographic proximity in satellite imagery with a triplet loss, while Pantazis et al. leverage spatiotemporal co-occurrence in wildlife monitoring. However, these methods are not directly applicable to histopathology, where (i) extremely high-resolution WSIs form grid-like spatial layouts, (ii) semantic class transitions can be subtle, and (iii) true negatives are difficult to define due to localized disease features.
@@
Figure 1: In standard SSL (blue), a patch is transformed to generate a positive pair. Our method (green) augments this by incorporating a spatially adjacent neighbor as a contextual pair. This leverages biological structure to enforce both transformation invariance and spatial coherence in representation learning.
In this paper, we propose a spatially coherent positive pair sampling strategy tailored for histopathology SSL. Rather than relying solely on augmentations, we incorporate the natural proximity of WSI patches to generate biologically meaningful positive pairs (Figure 1). Our approach is method-agnostic and can be integrated into contrastive or redundancy-reduction SSL frameworks.
@@
We evaluate our method across multiple datasets and architectures. Rather than aiming to surpass state-of-the-art performance, our goal is to present a biologically grounded method that enhances representation learning for histopathology, aligning with the growing interest in foundational models for medical imaging [21, 27, 25]. By leveraging spatial coherence between adjacent image patches, our approach constructs more meaningful positive pairs, leading to improved WSI classification performance over standard SSL sampling strategies.
2 Related Work
2.1 Deep Learning for Digital Histopathology
Deep learning (DL) approaches for whole-slide image (WSI) classification typically follow a multi-stage pipeline [36, 31, 3, 34]. A WSI is first segmented into smaller patches, often with non-tissue regions removed. These patches are then processed (either labeled at the patch level or treated as unlabeled instances) and their features aggregated into slide-level representations using attention-based pooling [9, 40, 44], feature maps [36, 31], or graphs [47, 3]. Finally, a downstream model performs slide-level classification.
However, despite the abundance of WSIs, two challenges limit supervised DL in pathology. First, generating high-quality annotations is time-intensive and requires expert pathologists, with risks of fatigue and inter-observer variability [6, 2]. Second, the gigapixel scale of WSIs makes end-to-end training computationally prohibitive. Furthermore, class imbalance and rare diseases exacerbate data scarcity for specific tasks [17, 12].
To overcome these barriers, several annotation-efficient approaches have been explored: transfer learning [42, 1, 37], synthetic data augmentation [20, 11, 39, 38], and annotation-free learning [19, 22, 33]. The latter, particularly self-supervised learning (SSL), has shown promise in extracting meaningful features without pixel-level labels.
2.2 Self-supervised Learning in Histopathology
Self-supervised learning (SSL) aims to learn visual representations from unlabeled data by solving proxy tasks. In histopathology, patch-level embeddings learned through SSL can be aggregated via multiple instance learning (MIL) to perform downstream WSI classification [44]. This paradigm reduces reliance on curated patch labels while achieving competitive performance.
Contrastive learning is a widely used SSL approach where a model learns to align representations of transformed views of the same input (positive pairs) while pushing apart other inputs (negatives) [8]. However, in WSIs, defining negatives is nontrivial: benign tissue dominates most slides, and spatially distant patches may still share morphology. Recent SSL frameworks such as BYOL, Barlow Twins, VICReg, and DINO avoid the need for negative pairs altogether [14, 46, 4, 7], focusing on redundancy reduction or view alignment instead.
Still, most SSL approaches rely heavily on data augmentations (e.g., cropping, color jitter, flipping) designed for natural images. These transformations may not adequately capture biological variations in tissue, especially structural continuity or local morphological patterns.
Foundation models for histopathology have recently emerged as a promising direction to scale representation learning across diverse WSI datasets and clinical tasks [27, 25, 21]. These models typically use large transformer-based architectures and are trained with extensive unlabeled data using objectives such as masked image modeling or contrastive learning. While this study is not designed to compete with or replace these large-scale models, it introduces a biologically motivated sampling strategy that can be integrated into existing SSL frameworks with minimal overhead. The incorporation of spatial context offers a way to improve patch-level representation quality, which could serve as a complementary enhancement to pretraining routines, especially in scenarios where data efficiency or morphological relevance is critical.
2.3 Positive Pair Sampling and Spatial Context
Recent work has shown that the choice of positive pairs can significantly affect SSL performance; sometimes more than the backbone architecture or loss function [30, 18, 24]. For example, Pantazis et al. leveraged spatial and temporal cues in wildlife camera imagery to form semantically meaningful pairs. Tile2Vec [18] introduced spatial sampling in satellite imagery, using geographic proximity and triplet loss to learn local structure-aware embeddings.
These works suggest that natural context (rather than synthetic augmentation) can provide rich supervisory signals. However, they have not been adapted to the unique structure of WSIs, which are high-resolution, grid-structured, and label-scarce.
In this work, we explore positive pair sampling based on spatial proximity within WSIs. Adjacent patches in a WSI typically exhibit strong morphological coherence, offering a biologically grounded signal for SSL. Unlike Tile2Vec, we do not assume distant patches are dissimilar: many WSIs, especially normal ones, contain spatially repetitive benign tissue. Also, we do not use a triplet loss or require negatives, aligning our approach with modern SSL practices.
Similar to our work is SimTriplet: both approaches exploit the spatial continuity of tissue morphology in WSIs to define biologically meaningful supervisory signals for self-supervised learning. Specifically, both methods construct positive pairs from spatially adjacent patches under the assumption that proximity implies semantic similarity in histopathology. However, our approach differs from SimTriplet in both architecture design and training objective. SimTriplet employs a triplet-based network that explicitly optimizes intra-sample and inter-sample similarities using a custom triplet structure and loss. In contrast, our method introduces a flexible contextual positive pair sampling strategy that does not require architectural changes or triplet loss formulations. Instead, we augment standard SSL pipelines by blending augmentation-based and context-based positive pairs through a weighted combination in the loss function (Equation 1). This modular design allows seamless integration into popular joint embedding SSL frameworks (e.g., BYOL, Barlow Twins, VICReg, DINOv2) without relying on handcrafted triplet structures or predictors. Additionally, we explicitly investigate the impact of spatial distance in contextual sampling, providing deeper insight into how spatial granularity influences representation learning in WSI data.
@@ 3 Methodology
3.1 Overview
We present a self-supervised learning (SSL) framework for whole-slide image (WSI) analysis that incorporates spatial context into positive pair selection. The pipeline consists of three main components: (1) patch extraction from WSIs, (2) SSL pretraining using contextual positive pair sampling, and (3) downstream evaluation via multiple instance learning (MIL) and linear probing. Our method is modular and compatible with a variety of joint embedding SSL approaches.
3.2 Contextual Positive Pair Sampling
We introduce a spatially-aware positive pair sampling strategy that leverages the local structural coherence of tissue morphology. The key idea is to incorporate spatially adjacent patches as contextual signals, augmenting standard augmentation-based sampling. An overview of this strategy is shown in Figure 1.
Let $U^{P}$ denote a set of $M$ pivot patches sampled from the WSIs in the training set. For each pivot patch $p^{i}_{j}\in U^{P}$, we identify a set of $K$ neighbor patches $U^{N}=\{p^{i}_{k}\}$ within the same WSI, constrained by a maximum Chebyshev distance. This metric is chosen due to its alignment with the grid structure of WSIs, where patches are spatially indexed.
During training, we construct two types of positive pairs:
  1. 1.
    Standard augmentation-based pair: Two augmented views of the same patch, $v_{1,j}=t_{1}(p^{i}_{j})$ and $v_{2,j}=t_{2}(p^{i}_{j})$, where $t_{1},t_{2}\sim T$ are sampled transformations.
  2. 2.
    Contextual pair: A neighbor patch $p^{i}_{k}\in U^{N}$ is randomly selected, and both $p^{i}_{j}$ and $p^{i}_{k}$ are augmented using the same transformation $t_{1}$, producing $v_{1,j}=t_{1}(p^{i}_{j})$ and $v_{1,k}=t_{1}(p^{i}_{k})$.
The combined loss function is defined as:
@@ 3.2 Contextual Positive Pair Sampling
where $L(\cdot,\cdot)$ denotes the loss function from the SSL framework in use (e.g., redundancy reduction for Barlow Twins or BYOL). The weight $\alpha\in[0,1]$ balances contextual similarity and transformation invariance.
This formulation avoids the need for negative pairs and is particularly well-suited for WSI data, where distant patches may still belong to the same class. The method is biologically motivated, computationally efficient, and compatible with both contrastive and non-contrastive SSL paradigms.
3.3 Integration with SSL Frameworks
Our sampling strategy is architecture-agnostic and integrates seamlessly with joint embedding SSL frameworks such as Barlow Twins, BYOL, VICReg, and DINOv2. During training, contextual and standard positive pairs are combined as defined in Equation 1, guiding the encoder to learn representations that are robust to both augmentations and local spatial variations. This design allows the approach to serve as a plug-and-play module within existing WSI pipelines, improving patch-level feature quality with minimal overhead.
@@ 4.1 Datasets
Private Datasets
We evaluated our method on two private datasets obtained from a collaborating medical institution: (1) 600 whole slide images (WSIs) of stomach tissue and (2) 600 WSIs of colorectal tissue. Each WSI comprises 2 to 4 sections from the same specimen. For each dataset, the slides were evenly categorized into normal/benign (N), dysplastic (D), and malignant (M) groups, with 200 slides per class. For each class, we randomly split the slides into 140 for training, 30 for validation, and 30 for testing. Patches of size 256$\times$256 pixels were extracted at 20$\times$ magnification using a sliding window approach implemented with the OpenSlide Python library [13].
Research use of these datasets was approved by the Institutional Review Boards of Institution A (Approval #InstA-IRB-2020–007) and Institution B (Approval #InstB-IRB-22-334, #InstB-2020-116). All WSIs were stained with hematoxylin and eosin (H&E), digitized in MIRAX format at a resolution of 200x using a Panoramic Flash 250 III scanner, and independently reviewed by two board-certified pathologists affiliated with Institution A. Patient data were fully anonymized prior to analysis.
@@ 4.2 Implementation Details
All experiments were conducted using PyTorch [32] on a Linux workstation equipped with a single NVIDIA GeForce RTX 3080 Ti GPU (32 GB). We evaluated four self-supervised learning (SSL) methods—Barlow Twins [46], BYOL [14], VICReg [4], and DINOv2 [29]—for training patch-level representation encoders. All methods used ResNet-18 [15] as the backbone, except DINOv2, which employed a Vision Transformer (ViT-Tiny) architecture.
For Barlow Twins, BYOL, and VICReg, we used the LARS optimizer [45] with a weight decay of $1\times 10^{-6}$. The learning rates were set to 0.2 for weights and 0.0048 for biases. These models were trained for 100 epochs with a batch size of 1024. For DINOv2, we adopted the AdamW optimizer with a cosine learning rate scheduler. The initial learning rate was set to 0.0005 with a weight decay of 0.04. The teacher’s temperature was linearly warmed up from 0.02 to 0.04 over 15 epochs. Due to computational constraints, DINOv2 was trained for 50 epochs using a batch size of 128.
Image Augmentation
To generate augmented views for self-supervised learning, we employed a diverse set of image transformations inspired by SimCLR [8], Kang et al. [21], and DINOv2 [29]. The augmentations included random horizontal and vertical flipping, random resized cropping, color jittering, and Gaussian blur. Additionally, for DINOv2, we adopted a multi-crop strategy consisting of two global crops and eight local crops. The global crops were generated using random resized cropping (scale range of 0.5–1.0) with bicubic interpolation, while the local crops were produced with smaller resized crop regions (scale range of 0.05–0.4). All crops were normalized using standard ImageNet mean and standard deviation statistics.
Slide-level Classification
To perform slide-level classification, we aggregated patch-level features using a Multiple Instance Learning (MIL) framework. In this setting, each WSI was treated as a bag of instances (i.e., extracted patches), and a single label was assigned to the entire bag rather than to individual patches. Specifically, we adopted the attention-based MIL (ABMIL) approach proposed in [16], which learns to assign attention weights to each patch feature. For each WSI, we first extracted patch embeddings using the pretrained encoders without any fine-tuning. These patch embeddings were then input to the ABMIL model to predict the slide-level label. The model was trained using the Adam optimizer with a learning rate of 0.001 and weight decay of $1\times 10^{-2}$. Training was conducted for 100 epochs.
Linear Probing
To evaluate the quality of the learned representations, we performed linear probing by training a shallow classifier on top of the frozen encoders obtained from the SSL methods. We sampled 5,000 patches per class for training and 2,500 patches per class for both validation and testing. The encoder weights were loaded from the corresponding SSL-pretrained checkpoints, and all encoder parameters were frozen during training. We evaluated both ResNet-18 and ViT-Tiny backbones. For ResNet-18, we used an SGD optimizer with an initial learning rate of 0.1 and applied cosine annealing as the learning rate scheduler. The linear classifier was trained for 50 epochs. For ViT-Tiny, we used the AdamW optimizer with a learning rate of 3e-4 and weight decay of 0.01, combined with a cosine scheduler and 5 warm-up steps over 50 training epochs. In both setups, the frozen encoder outputs were fed into a linear layer to perform tissue classification.
As a supervised baseline, we trained the same ResNet-18 and ViT-Tiny backbones from scratch using labeled patch data. The training configuration, including optimizers, learning rates, and schedulers, was identical to that used in the linear probing setup. However, unlike linear probing, the entire model (encoder and classifier) was trained end-to-end for 50 epochs.
We excluded the TCGA dataset from the linear probing and supervised training experiments, as it does not contain patch-level labels required for training or evaluating patch classifiers.
Performance Metrics
The performance of the models was evaluated using accuracy and the Area Under the Receiver Operating Characteristic Curve (AUCROC).
@@ 4.3 Contextual Positive Pair Sampling
Figure 2: Percentage of mismatches as a function of Chebyshev distance for three histopathology datasets (Stomach, Colon, and Breast). TCGA (Lung) was excluded because it does not contain patch-level annotations.